mmmaaa:

The p-type region of a silicon p-n junction is doped with 1016 boron atoms per cubic centimeter, and the n-type region is doped with 1018 phosphorus atoms per cubic centimeter. Assume a step p-n junction and that all doping atoms are ionized. The intrinsic carrier concentration in silicon at 300K is 1010cm−3. Calculate the width of the depletion region at 300 K (the relative permittivity of Si is r = 11.7).

CutieKittyGirl539:

Any choices

toga:

To calculate the width of the depletion region at 300K, we can use the following equation: W = √(2εrε0(Vbi + Vapplied)/q)(1/Nd + 1/Na) where: - W is the width of the depletion region - εr is the relative permittivity of Si - ε0 is the permittivity of free space - Vbi is the built-in potential of the p-n junction - Vapplied is the applied voltage across the junction - q is the elementary charge - Nd is the doping concentration of the n-type region - Na is the doping concentration of the p-type region At 300K, the built-in potential of the p-n junction can be calculated as follows: Vbi = kT/q * ln(Nd*Na/n_i^2) where: - k is Boltzmann's constant - T is the temperature in Kelvin - n_i is the intrinsic carrier concentration in silicon at 300K Plugging in the given values, we get: Vbi = 0.026V * ln(1018*1016/10^20) = 0.76V Assuming no applied voltage, Vapplied = 0, we can simplify the equation for the depletion region width as: W = √(2εrε0Vbi/q)(1/Nd + 1/Na) Plugging in the given values, we get: W = √(2*11.7*8.854*10^-12*0.76/1.6*10^-19)(1/1018 + 1/1016) = 1.14*10^-5 meters Therefore, the width of the depletion region at 300K is approximately 11.4 micrometers.